stock's returns have the following distribution:

Demand for the
Company's Products
Probability of This
Demand Occurring
Rate of Return If
This Demand Occurs
Weak 0.1 (26%)
Below average 0.2 (11)   
Average 0.3 17   
Above average 0.3 21   
Strong 0.1 64   
1.0

Assume the risk-free rate is 3%. Calculate the stock's expected return, standard deviation, coefficient of variation, and Sharpe ratio. Do not round intermediate calculations. Round your answers to two decimal places.

Stock's expected return: %

Standard deviation: %

Coefficient of variation:

Sharpe ratio:
    Answer #1

    Expected Return = [0.1(-0.26) + 0.2(-0.11) + 0.3(0.17) + 0.3(0.21) + 0.1(0.64)]

    Expected Return = 0.13

    Standard Deviation = [0.1(-0.26 - 0.13)2 + 0.2(-0.11 - 0.13)2 + 0.3(0.17 - 0.13)2 + 0.3(0.21 - 0.13)2 + 0.1(0.64 - 0.13)2]1/2

    Standard Deviation = 0.2348

    CV = 0.2348/0.13 = 1.81

    Sharpe Ratio = (0.13 -0.03)/0.2348

    Sharpe Ratio = 0.43